A project has estimated annual cash flows of $90,000 for 3 years and is estimated to cost $250,000. Assume a minimum acceptable rate of return of 10%.

Present Value of $1 at Compound Interest

Year 6% 10% 12%
1 0.943 0.909 0.893
2 0.890 0.826 0.797
3 0.840 0.751 0.712
4 0.792 0.683 0.636
5 0.747 0.621 0.567

Present Value of an Annuity of $1 at Compound Interest

Year 6% 10% 12%
1 0.943 0.909 0.893
2 1.833 1.736 1.690
3 2.673 2.487 2.402
4 3.465 3.170 3.037
5 4.212 3.791 3.605

Use the tables above.

a. Determine the net present value of the project. Enter negative values as negative numbers.
$

b. Determine the present value index. Round your answer to two decimal places.
Answer #1

a) Calculation of Net Present Value:

Present outflow = - $250,000

Present value of future inflows = 90,000 x Present value annuity factor @10%

= 90,000 x 2.487

= $ 223,830

So the net present value = 223830 - 250000 = - $26,170

b) Present Value index = Present value of future inflows ÷ Present Value of outflows

= 223830 / 250000 = 0.90